OpenStudy (anonymous):

What are the possible rational zeros of f(x) = 2x3 – 4x2 – 7x + 14 ?

OpenStudy (y2o2):

2x3 – 4x2 – 7x + 14 = 2x²(x-2)-7(x-2) = (x-2)(2x²-7) = 0 you can proceed...

OpenStudy (nenadmatematika):

just 2

OpenStudy (anonymous):

you can also use the rational root theorem..

OpenStudy (nenadmatematika):

the other zeros are +-sqrt(7/2) but they're not ratonal numbers

OpenStudy (anonymous):

possible rational roots would include all integer factors of 14 divided by all integer factors of the leading coefficient.

OpenStudy (anonymous):

so +-14/2,+-7/2,+-7/1,+-7/1

OpenStudy (anonymous):

that doesn't mean all are roots, it just means these are all of the possible roots.

OpenStudy (anonymous):

oops I forgot a couple ....+-2/2, +-2/1